Wallis Company manufactures only one product and uses a standard cost system. The company uses a predetermined plantwide overas follows The companys standard cost card for its only product (1) Standard Quantity (2) Standard Standard Price Cost (1) хRequired: 1. Compute all direct materials, direct labor, and fixed overhead variances for the year. 2. Record transactions aComplete this question by entering your answers the tabs below. Req 4 Req 1 Rea 2 and 3 Record transactions a through i for WComplete this question by entering your answers in the tabs below. Req 1 Req 4 Req 2 and 3 Prepare Wallis Companys income st
Answer #1

2 and 3

Wallis Company
Transaction Analysis
For the Year Ended 12/31/XX
(dollars in thousands)
Cash Raw Materials Work-in-Process Finished Goods PP&E (net) = Materials Price Variance Material Quantity Variance Labor Rate Variance Labor Efficiency Variance Fixed Overhead Budget Variance Fixed Overhead Volume Variance Retained Earnings
1/1 720.00 170.00 0.00 290.00 8700.00 = 0.00 0.00 0.00 0.00 0.00 0.00 9880.00
a. -6860.70 7022.40 = 161.70
b. -6551.20 5788.16 = -763.04
c. -3926.40 3712.80 = -736.20 522.60
d. -1342.00 2856.00 -1399.00 = 141.00 -26.00
e. -12356.96 12356.96 =
f. 15674.00 = 15674.00
g. -11967.56 = -11967.56
h. -2121.00 = -2121.00
i. = -161.70 763.04 736.20 -522.60 -141.00 26.00 -699.94
12/31 2143.90 641.20 0.00 679.40 7301.00 = 0.00 0.00 0.00 0.00 0.00 0.00 10765.50

Predetermined overhead rate = $2882000/288200 = $10 per direct labor hour

Fixed overhead applied = (95200 x 3 x $10)/1000 = $2856

4.

Wallis Company
Income Statement
For the Year Ended 12/31/XX
(dollars in thousands)
Sales revenue (92200 x $170/100) 15674
Cost of goods sold at standard (92200 x $129.80/1000) 11968
Total variance adjustments 699.94 12668
Gross profit 3007
Selling and administrative expenses 2121
Net operating income 885